OpenStudy (anonymous):

The equation of the normal line to the curve y = cube root of (x^2 - 1) at the point where x = 3 is

OpenStudy (anonymous):

First, you need to find the slope of the curve. You'll find it using derivative: dy/dx = (1/3)*[(x^2 - 1)^(-2/3)]*(2x) We'll calculate the slope at x=3 dy/dx = m = (1/3)*[(9 - 1)^(-2/3)]*(6) m = 2/8^(2/3) m = 2/4 m = 1/2 The slope of the normal line is m1 = -1/(1/2) = -2 Now, we'll calculate the y coordinate at x=3 y = cube root(9-1) = 2 The equation of normal line is: y - 2 = -2(x-3) y = -2x + 8